Accueil > Sites heberges > CFHTLS-Deep > 2004 runs

 

2004 runs

Par Genevieve Soucail - 24/04/2006

 

 Summary of 2004 runs

  • Summary of queue run 04BQ06 : 31 December - 21 January
    • Again a bad run with 10 nights lost to weather (ice, wind) or with seeing out of range for LS
    • Start of observations for D3.
    • D1 observed for 1.4 hours (validated), D2 for 6.6 hours, and D3 for 2.2 hours. Very small numbers indeed.
    • Poor filter balance as usual when the weather is bad. Not a single u* image taken during the run !!!

  • Summary of queue run 04BQ05 : 02 December - 21 December
    • New configuration of L3 (flipped) : significant improvement of the image quality accross the mosaic.
    • 4 nights entirely lost to weather (winter in Hawaii !) and 4 more nights with very bad seeing and unvalidated data.
    • D1 observed for 6.6 hours (validated) D2 for 6.9 hours, but with no u* data.

  • Observations of D4 are closed (Dec. 1, 2004). See the D4 web page for details

  • Summary of queue run 04BQ04 : 01 November - 22 November
    • A difficult start for this run, with some focus problems, identified later as due to a flip of L3. Surprisingly, significant improvement appeared in the IQ over a large part of the field. Waiting for further investigations, L3 was re-mounted in its nominal position for the rest of the run afer 4 nights of observations
    • 5 nights lost to bad weather over the run (fog, humidity, and bad seeing).
    • Good amount of data taken for the Deep, especially in u* on D1. About 10h of u* observations were put in high prioirity by the Steering Group for the run, from which more than 6h were observed and validated.
    • D1 observed for 16h with large amount of u*, D4 for 8.5 hours.
    • Start of the observations for D2 (reference field observed for 1.8h).

  • Summary of queue run 04BQ03 : 03 October - 22 October
    • A rather good run for the CFHTLS although a few nigths (or parts of) lost due to strong absorption and/or seeing out of specs. Plus one night lost to snow !
    • A rather satisfactoring amount of data taken for the Deep, with one sequence in u* on D1 but not on D4. Some sequences not validated for the Deep because the seeing was too high, especially in z’.
    • D4 observed for 6.8 hours and D1 observed for a total of 10.7 hours

  • Summary of queue run 04BQ02 : 04 September - 23 September
    • A difficult start for the run, with only half the mosaic working. Then it was rapidly resolved and the rest of the run was close to optimal : good and stable weather, seeing mostly good to outstanding (yes !!!). A lot of data taken for the CFHTLS.
    • However, due to a large accumulation of requested time close to RA=2h (CFHTLS Deep and Wide + highly ranked PI programs), no u*-band data were acquired on D1. The deficit in u* data on D1 starts to be an issue and we all hope that it will be partly absorbed during the next run, especially if weather cooperates again.
    • D4 observed for 12.1 hours and D1 observed for a total of 9.1 hours

  • Observations of D3 are closed (Sep. 1, 2004). See the D3 web page for details

  • Summary of queue run 04BQ01 : 06 August - 23 August
    • A difficult start for the run, with a lot of technical problems. Some of them being resolved during the first nights of the run (no validated data), the rest only partly affecting the quality of the data (see CFHTLS news pages for more details : some CCD failures, mostly on CCDs 00 and 11, some ice deposit during one night, creating additional structures on the images, and light contamination during the first half of the run, due to a vacuum gauge emitting inside the cryostat.
    • Good to excellent weather most of the time. Only one night lost to weather, not affecting the SNLS data. Good seeing, except for a few data.
    • End of observations for D3, with only r’ and i’ data (1.5h in total, affected by the CCD failures)
    • Routine observations for D4 and D1 observed for a total of 8.3h and 7.9h respectively. Only a few u* data taken on D4, not on D1.

  • Summary of queue run 04AQ07 : 07 July - 27 July
    • Good run. Most of the nights were photometric, seeing was good with some variability. A few technical problems with the telescope the last two nights of the run.
    • Limited observability for D3, with only 4.2 hours spent on the field, and again no u* observations. Whatever the reasons, this raises a serious problem for the science to be done presently.
    • More favorable situation for D4, observed for a total of 10.4h, with at least 92% acceptable for the Deep. Some complementary observations in u*, g’ and z’ taken out of the Supernovae programme for a better filter balance.
    • Beginning of observations for D1, with reference data in r’ and i’ (1.4h in total)

  • Summary of queue run 04AQ06 : 07 June - 26 June
    • A new rather successful run with no more than 2 nights in total lost to weather (humidity and cirrus), + some periods of bad seeing. One night lost because the MegaCam cooling system failed. It was repaired during day time.
    • Again, no u-band observations were done which is a pity for the Deep survey. Fortunately, the fraction of time acceptable for the Deep is higher.
    • D3 observed in g’r’i’z’ for a total of 8.5 hours, with 86% acceptable for the Deep. No u* observations
    • D4 observed in g’r’i’ for a total of 5.2 hours, with 96% acceptable for the Deep. No u* nor z’ observations. Only observations for the reference field were acquired.

  • Observations of D2 are closed (Jun. 1, 2004). See the D2 web page for details

  • Summary of queue run 04AQ04-5 : 07 May - 14 May and 19 May - 27 May
    • Rather successful run with no more than 2 nights lost to weather (humidity and fog), + some periods of very bad seeing. The run was splitted in two periods (04AQ04 and 04AQ05) with MegaPrime dismounted in the middle. A new scheme of observing blocks was tested for the Supernovae programme on D3, leading to a better balance of filters at final. Unfortunately, no u-band observations were done, mostly because dark time was over-subscribed by other PI programs.
    • D2 observed in g’r’i’z’ for a total of 4.6 hours, with 73% acceptable for the Deep. No u* observations. Last run for the period.
    • D3 observed in g’r’i’z’ for a total of 8.5 hours, with 86% acceptable for the Deep. No u* observations

  • Summary of queue run 04AQ03 : 12 Mar. - 31 Mar.
    • Finally not a bad run, after a difficult start. About 7 nights lost to weather (fog, humidity or very bad seeing), mostly at the beginning of the run. Variable seeing with significant improvements over the run.
    • D2 observed in g’r’i’ for a total of 3.7 hours, but 55% only acceptable for the Deep. No u* observations
    • D3 observed in u*g’r’i’z’ for a total of 8.7 hours, with 78% acceptable for the Deep.

  • Summary of queue run 04AQ02 : 12 Mar. - 31 Mar.
    • Again a bad run. 13 nights out of 19 lost to weather... Seeing larger than 0.9" 50% of the rest of the timee, mostly due to windshake.
    • No u* available, and g’ observations not usefull for the Deep because of the seeing range.
    • 6.3h of observations for D2 (g’r’i’z’) but only 40% can be acceptable for the Deep.
    • 5.8h of observations for D3 (g’r’i’z’) but only 40% can be acceptable for the Deep.

  • Summary of queue run 04AQ01 : 25 Feb. - 05 Mar.
    • No observations at all ! 22 nights lost : 15 due to technical problems with MegaCam and then 7 due to bad weather (fog, snow, storms...)
    • The only good new is that CCD03 has been repaired.

  • Observations of D1 are closed (Feb 1, 2004). See the D1 web page for details

  • Summary of queue run 03BQ07 : Jan. 11 - Jan. 30
    • Another terrible run with more than 2/3 of the time lost to weather. Observations only during the first half of the run, then wind, fog, snow ...
    • No u* nor z’ observations available.
    • Last observations of D1, only in r’ and i’ for a total of 3.2 hours.
    • D2 observed in g’, r’ and i’, in poor conditions (significant attenuation most of the time) for a total of 4.3 hours.
    • Beginning of observations for D3, only in r’ and i’ for a total of 2.7 hours.

Back to recent updates

 

 

[Page Précédente] [Dans la même rubrique] [Sommaire]

 

 

Site du CNRS Logo OMP
Logo UPS
Logo INSU-CNRS

Annuaire

Rechercher

Sur ce site



Bibliotheque

Laboratoire Astrophysique de Toulouse - Tarbes (UMR5572)

CNRS (Midi-Pyrénées)

Univ. Paul Sabatier

Liens utiles

Accueil Imprimer Contact mail Plan du site Crédits